Shawn Bixby borrowed $18,000 on a 150-day, 11% note. After 60 days, Shawn paid $2,100 on the note. On day 90, Shawn paid an additional $4,100. Use ordinary interest.

a. Determine the total interest use the U.S. Rule. (Do not round intermediate calculations. Round your answer to the nearest cent.)

b. Determine the ending balance due use the U.S. Rule. (Do not round intermediate calculations. Round your answer to the nearest cent.)
Answer #1
Workings:
Days for int. calculations Principal on which int.is calculated Interest Partial pmt. Principal on which int.is calculated
1 2(from 1) 3=Prev.6 4=3*11%*Col.2/360 5=Given 6=3+4-5
0 18000
0-60 60 18000 330.00 2100 16230.00
60-90 30 16230.00 148.78 4100 12278.78
90-150 60 12278.78 225.11 12503.89
150 703.89
Interest upto 1st partial payment after 60 days=
18000*11%*60/360=
330
Now, the principal is
18000+330-2100=
16230
Interest upto 2nd partial payment after 90 days=
16230*11%*30/360=
148.78
Now, the principal balance is
16230+148.78-4100=
12278.78
Interest at the final payment after 150 days=
12278.78*11%*60/360=
225.11
Now, the ending principal balance due is
12278.78+225.11=
12503.89
So, the ANSWERS are:
a.Total interest using the US rule
330+148.78+225.11=
703.89
ie.
704
&
b. The ending balance due use the U.S. Rule
12503.89
ie.
12504